# IterIS: Iterative Inference-Solving Alignment for LoRA Merging

### v1

@@ Abstract
Low-rank adaptations (LoRA) are widely used to fine-tune large models across various domains for specific downstream tasks. While task-specific LoRAs are often available, concerns about data privacy and intellectual property can restrict access to training data, limiting the acquisition of a multi-task model through gradient-based training. In response, LoRA merging presents an effective solution by combining multiple LoRAs into a unified adapter while maintaining data privacy. Prior works on LoRA merging primarily frame it as an optimization problem, yet these approaches face several limitations, including the rough assumption about input features utilized in optimization, massive sample requirements, and the unbalanced optimization objective. These limitations can significantly degrade performance. To address these, we propose a novel optimization-based method, named IterIS: 1) We formulate LoRA merging as an advanced optimization problem to mitigate the rough assumption. Additionally, we employ an iterative inference-solving framework in our algorithm. It can progressively refine the optimization objective for improved performance. 2) We introduce an efficient regularization term to reduce the need for massive sample requirements (requiring only 1-5% of the unlabeled samples compared to prior methods). 3) We utilize adaptive weights in the optimization objective to mitigate potential unbalances in LoRA merging process. Our method demonstrates significant improvements over multiple baselines and state-of-the-art methods in composing tasks for text-to-image diffusion, vision-language models, and large language models. Furthermore, our layer-wise algorithm can achieve convergence with minimal steps, ensuring efficiency in both memory and computation.
@@ 1 Introduction
Figure 3: Three key limitations of real-distribution-based merging methods and our improvements. In the scenario of combining the COLA [46] and MNLI [47] tasks in NLP, we compare the representative real-distribution-based merging method, RegMean [15], with our proposed method: (a) RegMean exhibits increasing discrepancies with deeper encoder layers, while IterIS can fully resolve discrepancies. The value of the “score” metric reflects the performance of the method. (b) Comparison of sample requirements and runtime proportions between our method and RegMean. (c) Proportional magnitudes of terms $T_{1}$ and $T_{2}$ in the optimization objective, demonstrating balanced values for our method and unbalance for RegMean.
@@ 2 Related Work
Parameter-Efficient Fine-Tuning (PEFT) [9] enables pre-trained models to adapt to downstream tasks with minimal additional parameters, reducing both computational overhead and storage needs. Recent PEFT methods include prompt tuning [14, 23, 4], adapter tuning [11, 49], and low-rank adaptation tuning (LoRA-tuning) [12]. Prompt tuning introduces a set of task-specific learnable tokens, while adapter tuning adds lightweight modules between layers to capture task-specific knowledge. Among various PEFT methods, LoRA-tuning has become one of the most widely adopted techniques. It has been widely applied for diffusion models and large language models for various downstream tasks [50, 12, 8, 51]. By incorporating low-rank decomposition matrices into select layers, LoRA captures task-specific nuances efficiently. Its memory and computational efficiency make it a powerful solution for fine-tuning large models.
@@ 3.1 Preliminaries
Building on the derivations above, we highlight three key limitations of exiting real-distribution-based merging methods: 1) Eq. (3) relies on rough assumption1, which assumes the input features for the unified adapter $\bm{W}$ is identical to $\bm{X}_{i}$ (i.e., in Eq. (3), $\bm{X}_{i}$ in $\bm{W}^{T}\bm{X}_{i}$ represents the approximated input features). Specifically, the input features for the first layer of the unified model are identical to those of the LoRA-tuned models when provided with the same input. However, as model depth increases, the discrepancies grow significantly, making this approximation rough and severely limiting performance. 2) To ensure the invertibility of $\sum_{i=1}^{N}\bm{X}_{i}\bm{X}_{i}^{T}$ in Eq. (4) and enhance robustness, they are usually accompanied by massive sample requirements. However, this is often impractical in real-world applications. 3) Since the weight of $\bm{W}_{i}$ in Eq. (4) is based on the inner product matrix $\bm{X}_{i}\bm{X}_{i}^{T}$, variations in the magnitudes of $\bm{X_{i}}$ can cause the solution to be overly influenced by the most prominent term. It potentially renders Eq. (3) an unbalanced optimization.
@@ 3.2 IterIS Algorithm
All $\tilde{\bm{X}}_{i}$ are updated iteratively, starting with $\tilde{\bm{X}}_{i}=\bm{X}_{i}$. At each iteration, IterIS has two steps: 1) Inference Step: By performing inference on the $i$-th task samples within the current unified model, we can extract $\tilde{\bm{X}}_{i}$. Here, $\tilde{\bm{X}}_{i}$ represents the input features for the unified adapter, 2) Solving Step: After updating the $\tilde{\bm{X}}_{i}$ in Eq. (6), we compute the solution $\bm{W}^{*}$ for each unified adapter. These unified adapters are then incorporated into the pre-trained model, resulting in a new unified model for the next iteration. By directly leveraging the input features for the unified adapter, our method mitigates the rough assumption1. Furthermore, through iterative refinement of the optimization objective, IterIS can achieve a more accurate representation of LoRA merging.
@@
Adaptive Weight Balancing. Setting uniform weights for $\lambda_{i}$ in Eq. (5) leads to the same unbalanced issues encountered in previous works. To mitigate this, we introduce adaptive weights for balancing, defined as follows:
$\small\lambda_{i}=\|\bm{W}_{i}\|_{F}^{2}\|\bm{W}_{i}^{T}\bm{X}_{i}\|_{F}^{-2}.$ (8)
This choice of $\lambda_{i}$ mitigates unbalances in the optimization process, preventing any terms from being overly prioritized.
Analysis of IterIS. By leveraging the directed acyclic graph structure in deep learning models, we derive an upper bound on the iteration count for our algorithm’s convergence2. Given an encoder with N-layer self-attention modules, where the Q, K, and V matrices are fine-tuned using LoRA, our algorithm can converge after $N-1$ iterations. To prevent overfitting, we cap the iterations at 20 in our experiments. Despite the increased number of iterations, IterIS converges with significantly fewer samples compared to previous methods, maintaining a low overall inference count and ensuring computational efficiency. Additionally, our method bypasses the computational demands of gradient-based training through closed-form solutions and layer-wise updates. Full process of IterIS is illustrated in the appendix, with the complete procedure detailed in Algorithm 1.
@@ 4.2 IterIS for Vision-Language Model
Experimental Setup. We employed IterIS to vision-language models for multi-style caption generation, utilizing BLIP-image-captioning-base [19]. Style-specific LoRAs were fine-tuned on BLIP. We used 50 prompt-image pairs per style-specific captioning task to derive input features. We set $\alpha=8\times 10^{-7}$ and 6 iterations. Our method was evaluated on the combination of two caption text styles. Specifically, we utilized the SentiCap dataset [24] to train style-specific LoRAs (“positive” and “negative”). We also conducted experiments with other caption style combinations, presented in the appendix.
@@
Metrics and Baselines. Following common practices [17], we evaluate the performance of the style caption generation using style accuracy[38], CIDEr [43], and BLEU 1-4 [28]. Apart from style accuracy, we report the average of all other metrics across test sets for both styles. Our method is benchmarked against linear merging and RegMean.[15].
Main Results. In Table 2, we present experimental results for the “positive” + “negative” style combination. We fine-tuned two specialized LoRA models: POS LoRA, focused on generating positive (POS) captions, and NEG LoRA, specialized in negative (NEG) captions. In single-style generation tests, each LoRA model performs well within its designated style but struggles to produce the other style, with POS LoRA achieving only 1.8% accuracy in NEG caption generation. Linear merging enables dual-style generation, though it results in limited style controllability, yielding only 52.2% accuracy for POS and 55.7% for NEG. RegMean improves style control, achieving 62.4% accuracy for POS and 69.2% for NEG captions. Our proposed method maintains the highest caption generation quality, surpassing RegMean by 20.7% in POS accuracy and 8.9% in NEG accuracy. Figure 6 shows examples of caption generation in POS and NEG styles utilizing our method. The unified adapters acquired by our method effectively control the generation style while maintaining both caption generation quality and diversity. For additional examples and detailed results, please refer to the appendix.
@@ 4.3 IterIS for Large Language Model
Metrics and Baselines. For in-domain task integration, we evaluate the performance of the unified model across all task combinations using the F1 score. For multi-task integration, Accuracy is used for all the tasks except COLA [46], which is assessed using the Matthews correlation coefficient (MCC) [25]. To align the MCC metric range with accuracy, we apply linear normalization to scale MCC values between 0 and 1. Our algorithm is benchmarked against linear merging and the state-of-the-art method, RegMean [15].
Main results. 1) For in-domain task integration, there are a total of 4 in-domain tasks, resulting in $C_{4}^{2}+C_{4}^{3}+C_{4}^{4}=11$ combinations. We present the average F1 scores across all 11 combinations for each task in the In-Domain section of Table 3. It can be observed that our IterIS significantly outperforms both the Linear and RegMean approaches, with improvements over RegMean of 3.6%, 2.1%, 0.3%, and 2.6% on all the in-domain tasks, respectively. 2) For Multi-Task integration, we tested all pairwise combinations, totaling $C_{7}^{2}=21$ combinations. We calculated the average metric for each GLUE sub-task across these 21 combinations, as shown in the Multi-Task section of Table 3. Our method consistently outperforms both Linear and RegMean, with the only exception being a slightly lower average score than linear merging on COLA. Notably, our method shows significant improvement over RegMean on MNLI, COLA, and MRPC, with increases of 1.7%, 1.4%, and 0.9%, respectively. Please refer to the appendix for detailed experimental results for each combination.
@@ 5 Conclusion
  • Workflow and Limitations (Section E): Finally, we illustrate the complete workflow of IterIS and conduct an analysis of the limitations of our approach.
@@ Appendix D More Results
More Results on V&L Model. As shown in Figure 6 of Section 4, we provide additional examples of positive and negative captions generated by our algorithm. A detailed version of Table 2 of Section 4 from the main text is presented in Table 11. Additionally, we conducted experiments on FlickrStyle10K [5], exploring combinations of two other distinct styles. The results are also included in Table 11. Due to the subjective nature of evaluating humor and romance, as well as the lack of the open test set, we didn’t quantify humor(romance)-specific metrics. As a result, aside from CIDEr score, the advantages of our algorithm for this style combination are less evident.
More Results on LLM. A detailed version of Table 3 in Section , as discussed in the main text, is provided for reference in this part. Table 12 summarizes the experimental results of 11 combinations for in-domain task integration, while Table 13 presents the results of 21 combinations for multi-task integration. Our proposed method outperforms both RegMean and linear merging in the majority of cases, highlighting its robustness and effectiveness.
Appendix E Other

### v2

@@ Abstract
Low-rank adaptations (LoRA) are widely used to fine-tune large models across various domains for specific downstream tasks. While task-specific LoRAs are often available, concerns about data privacy and intellectual property can restrict access to training data, limiting the acquisition of a multi-task model through gradient-based training. In response, LoRA merging presents an effective solution by combining multiple LoRAs into a unified adapter while maintaining data privacy. Prior works on LoRA merging primarily frame it as an optimization problem, yet these approaches face several limitations, including the rough assumption about input features utilized in optimization, massive sample requirements, and the unbalanced optimization objective. These limitations can significantly degrade performance. To address these, we propose a novel optimization-based method, named IterIS: 1) We formulate LoRA merging as an advanced optimization problem to mitigate the rough assumption. Additionally, we employ an iterative inference-solving framework in our algorithm. It can progressively refine the optimization objective for improved performance. 2) We introduce an efficient regularization term to reduce the need for massive sample requirements (requiring only 1-5% of the unlabeled samples compared to prior methods). 3) We utilize adaptive weights in the optimization objective to mitigate potential unbalances in LoRA merging process. Our method demonstrates significant improvements over multiple baselines and state-of-the-art methods in composing tasks for text-to-image diffusion, vision-language models, and large language models. Furthermore, our layer-wise algorithm can achieve convergence with minimal steps, ensuring efficiency in both memory and computation.00footnotetext: Work was done when Hongxu Chen and Zhen Wang were visiting students at HKUST. Long Chen is the corresponding author.
@@ 1 Introduction
Figure 3: Three key limitations of real-distribution-based merging methods and our improvements. In the scenario of combining the COLA [49] and MNLI [50] tasks in NLP, we compare the representative real-distribution-based merging method, RegMean [18], with our proposed method: (a) RegMean exhibits increasing discrepancies with deeper encoder layers, while IterIS can fully resolve discrepancies. The value of the “score” metric reflects the performance of the method. (b) Comparison of sample requirements and runtime proportions between our method and RegMean. (c) Proportional magnitudes of terms $T_{1}$ and $T_{2}$ in the optimization objective, demonstrating balanced values for our method and imbalance for RegMean.
@@ 2 Related Work
Parameter-Efficient Fine-Tuning (PEFT) [11] enables pre-trained models to adapt to downstream tasks with minimal additional parameters, reducing both computational overhead and storage needs. Recent PEFT methods include prompt tuning [17, 26, 6], adapter tuning [13, 53], memory-efficient tuning [3, 4], and low-rank adaptation tuning (LoRA-tuning) [14]. Prompt tuning introduces a set of task-specific learnable tokens, while adapter tuning adds lightweight modules between layers to capture task-specific knowledge. Among various PEFT methods, LoRA-tuning has become one of the most widely adopted techniques. It has been widely applied for diffusion models and large language models for various downstream tasks [54, 14, 10, 55]. By incorporating low-rank decomposition matrices into select layers, LoRA captures task-specific nuances efficiently. Its memory and computational efficiency make it a powerful solution for fine-tuning large models.
@@ 3.1 Preliminaries
Building on the derivations above, we highlight three key limitations of exiting real-distribution-based merging methods: 1) Eq. (3) relies on rough assumption1, which assumes the input features for the unified adapter $\bm{W}$ is identical to $\bm{X}_{i}$ (i.e., in Eq. (3), $\bm{X}_{i}$ in $\bm{W}^{T}\bm{X}_{i}$ represents the approximated input features). Specifically, the input features for the first layer of the unified model are identical to those of the LoRA-tuned models when provided with the same input. However, as model depth increases, the discrepancies grow significantly, making this approximation rough and limiting performance. 2) To ensure the invertibility of $\sum_{i=1}^{N}\bm{X}_{i}\bm{X}_{i}^{T}$ in Eq. (4) and enhance robustness, they are usually accompanied by massive sample requirements. However, this is often impractical in real-world applications. 3) Since the weight of $\bm{W}_{i}$ in Eq. (4) is based on the inner product matrix $\bm{X}_{i}\bm{X}_{i}^{T}$, variations in the magnitudes of $\bm{X_{i}}$ can cause the solution to be overly influenced by the most prominent term. It potentially renders Eq. (3) an imbalanced optimization.
@@ 3.2 IterIS Algorithm
All $\tilde{\bm{X}}_{i}$ are updated iteratively, starting with $\tilde{\bm{X}}_{i}=\bm{X}_{i}$. At each iteration, IterIS has two steps: 1) Inference Step: By performing inference on the $i$-th task samples within the current unified model, we can extract $\tilde{\bm{X}}_{i}$. Here, $\tilde{\bm{X}}_{i}$ represents the input features for the unified adapter, 2) Solving Step: After updating the $\tilde{\bm{X}}_{i}$ in Eq. (6), we compute the solution $\bm{W}^{*}$ for each unified adapter. These unified adapters are then incorporated into the pre-trained model, resulting in a new unified model for the next iteration. By directly leveraging the input features for the unified adapter, our method mitigates the rough assumption1. Furthermore, through iterative refinement of the optimization objective, IterIS can achieve a more accurate merging representation.
@@
Adaptive Weight Balancing. Setting uniform weights for $\lambda_{i}$ in Eq. (5) leads to the same imbalanced issues encountered in previous works. To mitigate this, we introduce adaptive weights for balancing, defined as follows:
$\small\lambda_{i}=\|\bm{W}_{i}\|_{F}^{2}\|\bm{W}_{i}^{T}\bm{X}_{i}\|_{F}^{-2}$ (8)
This choice of $\lambda_{i}$ mitigates imbalances in the optimization process, preventing any terms from being overly prioritized.
Analysis of IterIS. By leveraging the directed acyclic graph structure in deep learning models, we derive an upper bound on the iteration count for our algorithm’s convergence2. Given an encoder with J-layer self-attention modules, where the Q, K, and V matrices are fine-tuned using LoRA, our algorithm can converge after $J-1$ iterations. To prevent overfitting, we cap the iterations at 20 in our experiments. Despite the increased number of iterations, IterIS converges with significantly fewer samples compared to previous methods, maintaining a low overall inference count and ensuring computational efficiency. Additionally, our method bypasses the computational demands of gradient-based training through closed-form solutions and layer-wise updates. Efficiency analysis and full process of IterIS are illustrated in the appendix, with the complete procedure detailed in Algorithm 1.
@@ 4.1 IterIS for Text-to-Image Diffusion Model
Task arithmetic [16] (0.522, 0.557) 0.667 0.535 0.338 0.218 0.142
Ties merging [51] (0.468, 0.543) 0.739 0.519 0.330 0.214 0.141
Hyper-linear [52] (0.565, 0.539) 0.769 0.535 0.342 0.223 0.147
@@ 4.2 IterIS for Vision-Language Model
Experimental Setup. We employed IterIS to vision-language models for multi-style caption generation, utilizing BLIP-image-captioning-base [22]. Style-specific LoRAs were fine-tuned on BLIP. We used 50 prompt-image pairs per style-specific captioning task to derive input features. We set $\alpha=8\times 10^{-7}$ and 6 iterations. Our method was evaluated on the combination of two caption text styles. Specifically, we utilized the SentiCap dataset [27] to train style-specific LoRAs (“positive” and “negative”).
@@
Metrics and Baselines. Following common practices [20], we evaluate the performance of the style caption generation using style accuracy[41], CIDEr [46], and BLEU 1-4 [31]. Apart from style accuracy, we report the average of all other metrics across test sets for both styles. Our method is benchmarked against linear merging, RegMean [18], task arithmetic [16], ties merging [51], and hyper-linear [52].
Main Results. In Table 2, we present experimental results for the “positive” + “negative” style combination. We fine-tuned two specialized LoRA models: POS LoRA, focused on generating positive (POS) captions, and NEG LoRA, specialized in negative (NEG) captions. In single-style generation tests, each LoRA model performs well within its designated style but struggles to produce the other style, with POS LoRA achieving only 1.8% accuracy in NEG caption generation. Linear merging enables dual-style generation, though it results in limited style controllability, yielding only 52.2% accuracy for POS and 55.7% for NEG. RegMean improves style control, achieving 62.4% accuracy for POS and 69.2% for NEG captions. Our proposed method achieves the highest caption generation quality, outperforming all other methods and exceeding RegMean by 20.7% in POS accuracy and 8.9% in NEG accuracy. Figure 6 shows examples of caption generation in POS and NEG styles utilizing our method. The unified adapters acquired by our method effectively control the generation style while maintaining both caption generation quality and diversity. Additional examples and detailed results are in the appendix.
@@ 4.3 IterIS for Large Language Model
Metrics and Baselines. For in-domain task integration, we evaluate the performance of the unified model across all task combinations using the F1 score. For multi-task integration, Accuracy is used for all the tasks except COLA [49], which is assessed using the Matthews correlation coefficient (MCC) [28]. To align the MCC metric range with accuracy, we apply linear normalization to scale MCC values between 0 and 1. Our algorithm is benchmarked against linear merging and the state-of-the-art method, RegMean [18]. We additionally introduce three extra baselines for the multi-task integration subset, with results provided in the appendix.
Main results. 1) For in-domain task integration, there are a total of 4 in-domain tasks, resulting in $C_{4}^{2}+C_{4}^{3}+C_{4}^{4}=11$ combinations. We present the average F1 scores across all 11 combinations for each task in the In-Domain section of Table 3. It can be observed that our IterIS significantly outperforms both the Linear and RegMean approaches, with improvements over RegMean of 3.6%, 2.1%, 0.3%, and 2.6% on all the in-domain tasks, respectively. 2) For Multi-Task integration, we tested all pairwise combinations, totaling $C_{7}^{2}=21$ combinations. We calculated the average metric for each GLUE sub-task across these 21 combinations, as shown in the Multi-Task section of Table 3. Our method consistently outperforms both Linear and RegMean, with the only exception being a slightly lower average score than linear merging on COLA. Notably, our method shows significant improvement over RegMean on MNLI, COLA, and MRPC, with increases of 1.7%, 1.4%, and 0.9%, respectively. Please refer to the appendix for detailed results.
@@ 5 Conclusion
Acknowledgment. This work was supported by the Hong Kong SAR RGC Early Career Scheme (26208924), the National Natural Science Foundation of China Young Scholar Fund (62402408), Huawei Gift Fund, the HKUST Sports Science and Technology Research Grant (SSTRG24EG04), the School of the Gifted Young (USTC), and USTC Fellowship for Undergraduate Programs. In particular, Hongxu Chen (First Author) sincerely expresses his gratitude to his parents and his friends Yanghao Wang, Jiazhen Liu, and Jinyue Jiang for their unwavering support.
@@
  • Workflow and Analysis (Section E): Finally, we illustrate the complete workflow of IterIS and conduct an analysis of the efficiency and limitations of our approach.
@@ Appendix D More Results
Composition (MRPC, SST2) (SST2, QQP) (QQP, MRPC)
Task arithmetic (0.806, 0.937) (0.937, 0.841) (0.845, 0.811)
Ties merging (0.765, 0.943) (0.944, 0.823) (0.846, 0.806)
Hyper-linear (0.809, 0.933) (0.931, 0.849) (0.846, 0.801)
IterIS (Ours) (0.824, 0.951) (0.947, 0.855) (0.855, 0.816)
Table 14: Performance for multi-task integration for additional three baselines.
@@
More Results on V&L Model. As shown in Figure 6 of Section 4, we provide additional examples of positive and negative captions generated by our algorithm. A detailed version of Table 2 of Section 4 from the main text is presented in Table 11. Additionally, we conducted experiments on FlickrStyle10K [7], exploring combinations of two other distinct styles. The results are also included in Table 11. Due to the subjective nature of evaluating humor and romance, as well as the lack of the open test set, we didn’t quantify humor(romance)-specific metrics. As a result, aside from the CIDEr score, the advantages of our algorithm for this style combination are less evident.
More Results on LLM. A detailed version of Table 3 in Section, as discussed in the main text, is provided for reference in this part. Table 12 summarizes the experimental results of 11 combinations for in-domain task integration, while Table 13 presents the results of 21 combinations for multi-task integration. Our proposed method outperforms both RegMean and linear merging in the majority of cases, highlighting its robustness and effectiveness. Additionally, Table 14 shows the superiority of our method compared to three extra baselines in the multi-task integration subset.
Appendix E Others
@@
[…]
Detailed Analysis for the Regularization Term and Adaptive Weights. 1) The regularization term prevents irreversible situations in solutions, which often occur in few-sample cases. In such cases, at least one solution to the optimization objective in Eq.(5) renders it zero. This means the output features in each layer of the merged model are equal to those in the corresponding layer of each individual model. Due to the limited number of inference samples, the risk of overfitting to these samples is high. Notably, we can infer that the optimal solution to Eq.(1) is close to the linear solution based on Eq.(2). Applying identity matrices as regularization terms aligns the solution more closely with the linear solution, enhancing robustness. However, diagonal matrices in RegMean cannot achieve the same effect. 2) From a dimensional analysis perspective, the scale of the adaptive weights corresponds to that of $\|\bm{X}_{i}\|_{F}^{-2}$. This neutralizes the scale of $\bm{X}_{i}$ in the optimization objective ($\|\bm{W}_{i}^{T}\bm{X}_{i}-\bm{W}^{T}\tilde{\bm{X}}_{i}\|_{F}^{2}$), thereby ensuring that the optimization objective remains unbiased to the scale of $\bm{X}_{i}$.